Find the power dissipated in:

(i) an extension cord having a 0.06 Ω resistance, through which 5.00 A is flowing.

(ii) a cord utilizing thinner wire, with a resistance of 0.300 Ω.


1
Expert's answer
2021-01-22T11:26:44-0500

Power is the rate of flow of energy. The power can be calculated from the formula

"P = I^2R"

I = the amount of current flowing

R = the resistance through which current flowing

(i)

"P = (5.00)^2(0.06) \\

= 1.50 \;W"

(ii)

"P = (5.00)^2(0.300) \\

= 7.50 \;W"